# HaSPeR: An Image Repository for Hand Shadow Puppet Recognition

### v5

@@ Abstract
Hand shadow puppetry, benchmark, computer vision, dataset, deep learning, image classification, mobile application, ombromanie, shadowgraphy, silhouette, transfer learning
@@ 1 Introduction
Ombromanie, the ancient art of hand shadow puppetry, is a form of art that involves the mesmerizing interplay of light and shadow through the construction and manipulation of shadow figures or silhouettes on a surface, typically a screen or a wall, using one’s hands, body, or props [1]. The alias “cinema in silhouette”111 https://en.wikipedia.org/wiki/Shadowgraphy_(performing_art) is sometimes used to refer to this proto-cinematic medium of entertainment. Its working principle is very straightforward—the puppeteer adeptly positions their hands between a radiant light source and a translucent screen, consequently conjuring shadows and silhouettes that emulate different creatures, as shown in Figure 3.
@@
In tandem with this motivation, this work introduces a seminal addition to the realm of data resources, HaSPeR (Hand Shadow Puppet Image Repository), a methodically curated novel image dataset of hand shadow puppets. The dataset comprises an assemblage of 15,000 samples, that we painstakingly source and verify from 68 professional shadowgraphist clips and 90 amateur shadowgraphist clips. We label and categorize the images with utmost precision to elicit robustness in the image classification models that will undergo training with these images. The samples in HaSPeR are diverse in nature since the source clips are recorded in a plethora of different poses, orientations, and background lighting conditions of the translucent screen. We conduct a detailed analysis of HaSPeR’s statistical characteristics. We also employ a variety of state-of-the-art (SOTA) pretrained image classification models to establish a performance benchmark for validating the integrity of the dataset. Additionally, we conduct a thorough evaluation of several facets of the ace ResNet34 model, including its feature representations, interpretability, explainability, and classification errors that it encounters. In an effort to assess the potential of digitized ombromanie teaching tools, we create a simple and lightweight prototype Android application using Flutter for classifying hand shadow puppet images from the phone’s camera feed. We posit that our dataset possesses the potential to offer a wealth of opportunities for exploration and analysis into the artistic domain of hand shadow puppetry.
@@ 3.2 Extracting Samples
To mitigate the presence of excessively similar and redundant image samples, we extract frames from these clips at reasonable intervals of $k$ after downsampling the clips to a resolution of $1438\times 1080$. The values of $k$ are judiciously chosen for the clips of each class, and every $k$th frame is selected as a candidate image sample (e.g., with $k\approx 180$, $200$, $220$ for a 60 FPS clip). Table 1 encapsulates some essential statistical information related to HaSPeR and provides a superficial overview of the dataset.
@@ 4.1 Background Variance
Figure 5: Samples with different silhouette properties.
@@ 5.3 Error Analysis
The confusion matrix for the ResNet34 model on our dataset, presented in Figure 12, reveals that the ‘Crab’ class exhibits the highest count of misclassifications. One obvious reason for this is the somewhat significant inter-class similarity among the ‘Bird’, ‘Moose’, ‘Rabbit’, and ‘Crab’ classes. Most of the misclassified samples are from visually similar classes. We can posit that navigating the intricacies of visually similar classes poses a significant challenge in this image classification task, as evident from the other pale-red entries of the confusion matrix in Figure 12. Even to the keen human eye, distinguishing between these classes may be perplexing, as they share common visual features, shapes, or color patterns that result in a high degree of resemblance. We examine various aspects, such as the distinctive features or characteristics that might have led to confusion and the degree of similarity between the misclassified classes. Figure 11(a) and 11(f) show the confusion between a ‘Crab’ sample and a ‘Rabbit’ sample which look visually quite similar. The same holds for one of the six ‘Moose’ samples that are misclassified as ‘Bird’ samples by the ResNet34 model, as depicted in Figure 11(b) and 11(g). We observe that misclassifications of this type occur when images belonging to different, but visually akin categories, are erroneously assigned to the wrong class. Another reason for misclassifications is the ambiguity of shape present in mid-action frames. For instance, the ‘Bird’ sample in Figure 11(h) is a transition frame between two successive wing flaps. However, due to the presence of this ambiguous sample in the training set of the ‘Bird’ class, the ResNet34 ends up misclassifying the ‘Panther’ sample in Figure 11(c) as a ‘Bird’ sample. The misclassification portrayed by the pair of Figure 11(d) and 11(i) is due to the combination of poor lighting and ineptitude of the amateur child puppeteer in creating ‘Llama’ shadows. The model confuses the ear protrusions of the ‘Llama’ sample to be the tentacular eyes of a ‘Snail’ sample. For Figure 11(e), we can pontificate the misclassification reasons to be the overlapping of the fingers and the distorted angle at which the sample was captured. The ‘Snail’ sample thereby gets wrongly classified as a ‘Bird’ sample due to the existence of the analogous sample portrayed in Figure 11(j). The green entries along the diagonal of the confusion matrix in Figure 12 indicate the reasonably good classwise prediction performance of the ResNet34 model, which is, to some extent, due to the perfectly balanced sample distribution in HaSPeR.
@@ 6 Conclusion and Future Work
In this paper, we explore an intriguing subject matter, hand shadow puppetry, in the realm of computer vision. We introduce HaSPeR, a novel dataset with a sizable collection of 15,000 hand shadow puppet images distributed across 15 classes, taken from both expert and amateur puppetry performance clips. We fine-tune 31 pretrained image classification models on HaSPeR to establish a benchmark for the dataset. To explain and analyze the performance of the most erudite model, ResNet34, in comparison with other baseline models, we visually manifest its feature space using the $t$-SNE dimensionality reduction technique. We also perform thorough qualitative and error analyses for the ResNet34 model. We envisage the possibility of developing applications for imparting the art of shadowgraphy, via mobile and embedded devices. We claim that this work is novel and significant since it is the first publicly available dataset and study on image classification benchmarking that focuses only on ombromanie. There are many avenues in our work that warrant further investigation. We hope to reconcile those desiderata by enriching our dataset with numerous permutations of arm positions and finger movements, preferably by employing more skilled individuals with varying palm and wrist structures, thereby creating more diverse silhouettes. We also plan to experiment with a gesture detection technology such as MediaPipe [^11] MediaPipe — https://developers.google.com/mediapipe or Microsoft Kinect [^12] Kinect for Windows — https://learn.microsoft.com/en-us/windows/apps/design/devices/kinect-for-windows for leveraging depth coordinates of hand landmarks [62], and assess their efficacy in classifying hand shadow puppets.

### v6

@@ Abstract
Hand shadow puppetry, art digitization, benchmark, computer vision, dataset curation, deep learning, image classification, mobile application, ombromanie, optical flow estimation, shadowgraphy, silhouette classification, transfer learning
@@ 1 Introduction
Ombromanie, the ancient art of hand shadow puppetry, is a form of art that involves the mesmerizing interplay of light and shadow through the construction and manipulation of shadow figures or silhouettes on a surface, typically a screen or a wall, using one’s hands, body, or props [1, 2]. The alias “cinema in silhouette”111 https://en.wikipedia.org/wiki/Shadowgraphy_(performing_art) is sometimes used to refer to this proto-cinematic medium of entertainment. Its working principle is very straightforward—the puppeteer adeptly positions their hands between a radiant light source and a translucent screen, consequently conjuring shadows and silhouettes that emulate different creatures, as shown in Figure 3.
@@
In tandem with this motivation, this work introduces a seminal addition to the realm of data resources, HaSPeR (Hand Shadow Puppet Image Repository), a methodically curated novel image dataset of hand shadow puppets. The dataset comprises an assemblage of 15,000 samples, that we painstakingly source and verify from 68 professional shadowgraphist clips and 90 amateur shadowgraphist clips. We label and categorize the images with utmost precision to elicit robustness in the image classification models that will undergo training with these images. The samples in HaSPeR are diverse in nature since the source clips are recorded in a plethora of different poses, orientations, and background lighting conditions of the translucent screen. We also inculcate silhouette motion diversity via optical flow estimation [7] in the frame extraction process. We conduct a detailed analysis of HaSPeR’s statistical characteristics. We also employ a variety of state-of-the-art (SOTA) pretrained image classification models to establish a performance benchmark for validating the integrity of the dataset. Additionally, we conduct a thorough evaluation of several facets of the ace ResNet34 model, including its feature representations, interpretability, explainability, and classification errors that it encounters. In an effort to assess the potential of digitized ombromanie teaching tools, we create a simple and lightweight prototype Android application using Flutter for classifying hand shadow puppet images from the phone’s camera feed. We posit that our dataset possesses the potential to offer a wealth of opportunities for exploration and analysis into the artistic domain of hand shadow puppetry.
@@ 3.2 Extracting Samples
To mitigate the presence of excessively similar and redundant image samples, we extract frames from these clips at reasonable intervals of $k$ after downsampling the clips to a resolution of $1438\times 1080$. The values of $k$ are judiciously chosen for the clips of each class, and every $k$th frame is selected as a candidate image sample (e.g., with $k\approx 180$, $200$, $220$ for a 60 FPS clip). From this sequence of extracted candidate frames, we prioritize those exhibiting significant motion. To this end, we estimate the motion vector field by calculating the optical flow between the consecutive $t$th frame and the $(t+k)$th frame. The magnitude of this motion is quantified by the mean L2 norm of the resulting flow field. We retain the frame pairs with an average flow magnitude surpassing a certain requisite threshold $\tau$, thereby ensuring the inclusion of dynamically distinct frames. We synergistically amalgamate two optical flow estimation methods: the Lucas–Kanade (LK) method [33] and the Total Variation L1 Regularization (TV-$L^{1}$) method [34]. The undergirding assumption beneath the LK method is brightness constancy and spatial coherence of the flow in a local neighbourhood of the pixel (say, the patch $W$) under consideration. It employs a multi-scale gradient descent optimization approach for the constraint equation shown in Equation 1.
$I_{x}\cdot u+I_{y}\cdot v+I_{t}=0$ (1)
where, $I_{x}=\frac{\partial I}{\partial x}$ and $I_{y}=\frac{\partial I}{\partial y}$ are the spatial gradients of the image intensity $I$, and $I_{t}=\frac{\partial I}{\partial t}$ is the temporal gradient. $u$ and $v$ are the horizontal and vertical components of the optical flow vector, respectively. The motion is then estimated by iteratively minimizing the cost function in Equation 2 at increasingly granular image resolutions, from coarse to fine.
$\begin{bmatrix}u\\ v\end{bmatrix}_{\text{LK}}=\operatorname*{arg\,min}_{u,v}\sum_{\{i,j\}\in W}\left[I_{x}(i,j)\cdot u_{ij}+I_{y}(i,j)\cdot v_{ij}+I_{t}(i,j)\right]^{2}$ (2)
which can be then solved with the closed-form solution in Equation 3.
$\begin{bmatrix}u\\ v\end{bmatrix}_{\text{LK}}=\begin{bmatrix}\sum_{\{i,j\}\in W}I_{x}^{2}&\sum_{\{i,j\}\in W}I_{x}I_{y}\\ \sum_{\{i,j\}\in W}I_{x}I_{y}&\sum_{\{i,j\}\in W}I_{y}^{2}\end{bmatrix}^{-1}\begin{bmatrix}-\sum_{\{i,j\}\in W}I_{x}I_{t}\\ -\sum_{\{i,j\}\in W}I_{y}I_{t}\end{bmatrix}$ (3)
As evident from Figure 4(c), the LK method can track the edge movement of the homogeneous silhouette patches, given the slight texture offered by the penumbral region of the shadow. However, as it is limited by local window constraints, it fails to capture the global motion of the shadow puppet. To ameliorate this issue, we resort to the TV-$L^{1}$ method, which is a variational method that minimizes the total variation of the flow field, subject to the L1 norm of the data fidelity term, which together form the energy function $E$ in Equation 4.
$\begin{bmatrix}u\\ v\end{bmatrix}_{\text{TV-}L^{1}}$ $=\operatorname*{arg\,min}_{u,v}E(u,v)$ (4)
[…]
$\begin{bmatrix}u\\ v\end{bmatrix}^{*}$ $=\operatorname*{arg\,max}_{u,v}\left(\left\lVert\begin{bmatrix}u\\ v\end{bmatrix}_{\text{LK}}\right\rVert_{2},\left\lVert\begin{bmatrix}u\\ v\end{bmatrix}_{\text{TV-}L^{1}}\right\rVert_{2}\right)$ (5)
Figure 4: Optical flow estimation of contiguous candidate frames from the ‘Horse’ class.
Then from this maximum combination flow field of resolution $M\times N$, we compute the average L2 norm $\overline{V}$ using Equation 6.
[…]
If $\overline{V}>\tau$, we retain the corresponding frame pair as candidate samples, otherwise we continue the process with the $(t+k)$th and the $(t+2k)$th frames. Table 1 encapsulates some essential statistical information related to HaSPeR and provides a superficial overview of the dataset.
@@ 4.2 Nature of the Silhouettes
Figure 6: Samples with different silhouette properties.
@@ 5.3 Error Analysis
The confusion matrix for the ResNet34 model on our dataset, presented in Figure 12, reveals that the ‘Crab’ class exhibits the highest count of misclassifications. One obvious reason for this is the somewhat significant inter-class similarity among the ‘Bird’, ‘Moose’, ‘Rabbit’, and ‘Crab’ classes. Most of the misclassified samples are from visually similar classes. We can posit that navigating the intricacies of visually similar classes poses a significant challenge in this image classification task, as evident from the other pale-red entries of the confusion matrix in Figure 12. Even to the keen human eye, distinguishing between these classes may be perplexing, as they share common visual features, shapes, or color patterns that result in a high degree of resemblance. We examine various aspects, such as the distinctive features or characteristics that might have led to confusion and the degree of similarity between the misclassified classes. Figure 13(a) and 13(f) show the confusion between a ‘Crab’ sample and a ‘Rabbit’ sample which look visually quite similar. The same holds for one of the six ‘Moose’ samples that are misclassified as ‘Bird’ samples by the ResNet34 model, as depicted in Figure 13(b) and 13(g). We observe that misclassifications of this type occur when images belonging to different, but visually akin categories, are erroneously assigned to the wrong class. Another reason for misclassifications is the ambiguity of shape present in mid-action frames. For instance, the ‘Bird’ sample in Figure 13(h) is a transition frame between two successive wing flaps. However, due to the presence of this ambiguous sample in the training set of the ‘Bird’ class, the ResNet34 ends up misclassifying the ‘Panther’ sample in Figure 13(c) as a ‘Bird’ sample. The misclassification portrayed by the pair of Figure 13(d) and 13(i) is due to the combination of poor lighting and ineptitude of the amateur child puppeteer in creating ‘Llama’ shadows. The model confuses the ear protrusions of the ‘Llama’ sample to be the tentacular eyes of a ‘Snail’ sample. For Figure 13(e), we can pontificate the misclassification reasons to be the overlapping of the fingers and the distorted angle at which the sample was captured. The ‘Snail’ sample thereby gets wrongly classified as a ‘Bird’ sample due to the existence of the analogous sample portrayed in Figure 13(j). The green entries along the diagonal of the confusion matrix in Figure 12 indicate the reasonably good classwise prediction performance of the ResNet34 model, which is, to some extent, due to the perfectly balanced sample distribution in HaSPeR.
@@ 6 Conclusion and Future Work
In this paper, we explore an intriguing subject matter, hand shadow puppetry, in the realm of computer vision. We introduce HaSPeR, a novel dataset with a sizable collection of 15,000 hand shadow puppet images distributed across 15 classes, taken from both expert and amateur puppetry performance clips, by dint of a sophisticated frame extraction process involving optical flow estimation. We fine-tune 31 pretrained image classification models on HaSPeR to establish a benchmark for the dataset. To explain and analyze the performance of the most erudite model, ResNet34, in comparison with other baseline models, we visually manifest its feature space using the $t$-SNE dimensionality reduction technique. We also perform thorough qualitative and error analyses for the ResNet34 model. We envisage the possibility of developing applications for imparting the art of shadowgraphy, via mobile and embedded devices. We claim that this work is novel and significant since it is the first publicly available dataset and study on image classification benchmarking that focuses only on ombromanie. There are many avenues in our work that warrant further investigation. We hope to reconcile those desiderata by enriching our dataset with numerous permutations of arm positions and finger movements, preferably by employing more skilled individuals with varying palm and wrist structures, thereby creating more diverse silhouettes. We also plan to experiment with a gesture detection technology such as MediaPipe [^11] MediaPipe — https://developers.google.com/mediapipe or Microsoft Kinect [^12] Kinect for Windows — https://learn.microsoft.com/en-us/windows/apps/design/devices/kinect-for-windows for leveraging depth coordinates of hand landmarks [66], and assess their efficacy in classifying hand shadow puppets.
@@ 7 Acknowledgments
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
Data Availability
The codes and datasets generated, written, and/or analyzed during the research project are available in the HaSPeR GitHub repository (https://github.com/Starscream-11813/HaSPeR) and in the HaSPeR Hugging Face repository (https://huggingface.co/datasets/Starscream-11813/HaSPeR).
Funding Sources
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors. All associated costs were self-funded by the authors. Access to computing resources was provided by the Systems and Software Lab (SSL) of the Islamic University of Technology (IUT).